OpenStudy (anonymous):

solving systems of equations solve algebraically, y-x^2=3x y=x-1

OpenStudy (anonymous):

so what is ur question

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

\[y=x^2+3x\] \[y=x-1\] same y, so \[x^2+3x=x-1\] or \[x^2+2x+1=0\] \[(x+1)^2=0\] etc

OpenStudy (anonymous):

thats not what the answer is meant to be. I think either substitution or elimination is supposed to be used. the answer should be two numbers

OpenStudy (anonymous):

that is algebraic

OpenStudy (anonymous):

and there are not two solutions, \((-1,-2)\) is the only solution

OpenStudy (anonymous):

it is two numbers, namely the x and the y, but it is one solution

OpenStudy (anonymous):

where did those 2 numbers come from?

OpenStudy (anonymous):

what is x if \((x+1)^2=0\) ?

OpenStudy (anonymous):

-1 and 1 not 2?